# Tool-Aided Evolutionary LLM for Generative Policy Toward Efficient Resource Management in Wireless Federated Learning

### v1

@@ Abstract
Federated Learning (FL) enables distributed model training across edge devices in a privacy-friendly manner. However, its efficiency heavily depends on effective device selection and high-dimensional resource allocation in dynamic and heterogeneous wireless environments. Conventional methods demand a confluence of domain-specific expertise, extensive hyperparameter tuning, and/or heavy interaction cost. This paper proposes a Tool-aided Evolutionary Large Language Model (T-ELLM) framework to generate a qualified policy for device selection in a wireless FL environment. Unlike conventional optimization methods, T-ELLM leverages natural language-based scenario prompts to enhance generalization across varying network conditions. The framework decouples the joint optimization problem mathematically, enabling tractable learning of device selection policies while delegating resource allocation to convex optimization tools. To improve adaptability, T-ELLM integrates a sample-efficient, model-based virtual learning environment that captures the relationship between device selection and learning performance, facilitating subsequent group relative policy optimization. This concerted approach reduces reliance on real-world interactions, minimizing communication overhead while maintaining high-fidelity decision-making. Theoretical analysis proves that the discrepancy between virtual and real environments is bounded, ensuring the advantage function learned in the virtual environment maintains a provably small deviation from real-world conditions. Experimental results demonstrate that T-ELLM outperforms benchmark methods in energy efficiency and exhibits robust adaptability to environmental changes.
Index Terms:
Large language model, generative policy, wireless federated learning, and resource management.
@@ I Introduction
Next-Generation (xG) wireless communication systems are envisioned to support various intelligent applications and services[1, 2], empowered by the exponential growth of wireless edge devices, such as mobile phones and sensors. As a prominent paradigm[3, 4, 5], Federated Learning (FL) emerges by facilitating collaborative model training across decentralized devices while maintaining data locality in a privacy-friendly manner. Nevertheless, the deployment of FL in wireless networks faces severe challenges, arising from the underlying heterogeneous computation and communication capabilities across devices [6, 7, 8, 9]. Correspondingly, a proper resource management policy for FL, which selects the participating devices and calibrates the utilized communications and computing resources, will need to be developed in dynamic wireless environments [10, 11, 12]. Typically, such a problem results in a high-dimensional optimization problem that can be partially solved by heuristic solutions [13, 14] or Reinforcement Learning (RL)-based approaches [15, 16, 17, 18, 19]. However, heuristic solutions often demand a confluence of domain-specific expertise and extensive tuning before adapting to unseen scenarios, while RL-based approaches require heavy interactions with the environment and become sluggish for large system dimensions and changing system dynamics [20, 21, 22]. Therefore, there is a strong incentive to find alternative solutions[23, 24].
@@ I-A Related Works
To implement FL over wireless networks, in each training round, edge devices upload the locally trained updates to a centralized server via wireless links, in exchange for aggregated models. Through several training rounds, the performance of the eventually learned global model is primarily impacted by resource and data heterogeneity. For example, disparities in computational and communication capabilities across devices lead to imbalanced time and energy consumption, while varying dataset sizes and non-Independent and Identically Distributed (non-IID) data distributions can result in biased model updates. Both would degrade the learning effectiveness [7, 8, 9]. Additionally, the deployment of FL encounters constraints due to the energy and time [14] budget [16, 25, 26]. For instance, the authors in [27, 28] formulate the joint learning and communication problem as the minimization of total energy consumption under latency constraints and provide computationally efficient closed-form solutions to optimize critical resources, including bandwidth allocation, CPU frequency, and transmission power. But, full device participation is bluntly assumed in [27, 28]. However, mobilizing the participation of a subset of qualified edge devices in each training round can contribute to improving learning efficiency. For example, biasing client selection towards clients with higher local loss [29] proves to yield faster convergence than the random selection in Federated Averaging (FedAvg) [3]. Also, a timer can be heuristically set to avoid the participation of straggler devices [14]. In [13], a guided participant selection scheme improves FL performance by jointly optimizing system and data utility through an epsilon-greedy algorithm. Moreover, a joint device scheduling and bandwidth allocation policy, which maximizes model accuracy within a fixed training time budget by a hybrid greedy-optimization approach, is proposed in [30], while energy constraints are neglected therein. More importantly, adapting these heuristics to unexpected scenarios, such as tuning hyperparameters to approximate FL performance limits[14, 30] or setting appropriate timers[13] when system bandwidth or device resources (e.g., maximum transmission power, dataset size) change, typically demands iterative adjustments and domain-specific expertise.
@@
RL provides an alternative methodology to solve wireless FL problems. Through continuously interacting with the environment, RL agents evaluate candidate policies in the wireless FL environment and gradually learn appropriate real-time policies. For example, [18] develops a Deep RL (DRL) method to select devices to upload their local model for aggregation, to maximize validation accuracy. Nevertheless, it overlooks energy and time constraints and assumes all devices participate in local training. Similarly, [17] adopts a Q-learning-based method to identify near-optimal participating devices. Furthermore, to ameliorate the impact of high dimensionality on single-agent RL with large-scale devices, [15] introduces a Multi-Agent Reinforcement Learning (MARL) framework that jointly optimizes model accuracy, processing latency, and communication efficiency. On the other hand, [19] improves the energy efficiency of FL by managing the CPU-cycle frequency of mobile devices based on DRL. To further improve the efficiency of FL, the works in [31, 32, 33] leverage DRL-based approaches to jointly optimize the device selection and bandwidth allocation. However, the above DRL-based methods only optimize partial parameters and rely on extensive training in a given environment. Therefore, the changes in environment or data often imply cumbersome retraining or even a fresh redesign of RL. Meanwhile, direct interaction with the environment incurs substantial communication overhead. Although imitation learning [21] can acquire a policy by learning from the optimal trajectories to reduce interaction overhead, there are no such expert systems strictly in this FL scenario. Besides, the training difficulty, which increases exponentially along with the potentially involved devices and managed resources, limits the practicability [20, 22].
@@
The remarkable capabilities of transformer-based Large Language Models (LLMs) [34, 35, 36, 37, 38] demonstrate the feasibility of a single general-purpose model, trained on extensive text corpora, for diverse, generalizable task accomplishment [39]. Capitalizing on the success of LLMs, we resort to an LLM-driven textual description approach to leverage its inherent flexibility and generalization. Leveraging fine-tuned LLMs for FL optimization starkly contrasts with those efforts, which focus on enhancing the training capabilities of LLMs in wireless FL [40, 41]. Notably, while LLMs exhibit strong linguistic abilities, the fine-tuning performance is still limited by the scope of the utilized training dataset [42]. Therefore, due to the non-availability of decision-making expert data for FL efficiency, fine-tuned LLMs cannot reliably interpret wireless FL data and generate effective decisions. On the other hand, training an LLM to map observations to actions requires vast real-world datasets, which are time-consuming and costly to collect, hindering scalability [43]. Moreover, though reasoning-enhanced LLMs use long token sequences to improve computational efficiency[44], this approach is infeasible in wireless FL systems due to time constraints. Therefore, due to the domain-specific expertise and computationally intensive optimization operations, extra effort is still needed to make LLMs qualified for decision-making in wireless FL.
@@ I-B Contributions
In this paper, we propose a Tool-aided Evolutionary LLM (T-ELLM) approach to realize efficient device selection and resource allocation in wireless FL. Specifically, T-ELLM adopts a natural language-based scenario prompt, where linguistic flexibility helps to improve the generalization capability[45]. Afterward, on top of a mathematics-established problem decoupling, T-ELLM capably yields device selection and resource management results by answering the scenario query and invoking convex optimization tools, respectively. Furthermore, to better tackle possible environmental changes, alongside available convex optimization tools, T-ELLM builds a sample-efficient model-based virtual learning environment, which characterizes the relationship between learning effectiveness and device selection, and enables the evolution of the inner LLM at a reduced communication cost. The main contributions of this paper are summarized as follows:
  • We consider a high-dimensional decision-making problem in wireless FL scenarios, and adopt a tool-based evolutionary LLM approach, which is referred to as T-ELLM. In particular, we fine-tune an LLM, which can respond to the linguistic description of wireless FL scenarios with an appropriate output of device selection.
@@ I-C Paper Structure
The remainder of this paper is organized as follows. We introduce the system model and problem formulation in Section II. Afterward, Section III presents how to decouple the optimization problem and explains the rationale behind it. In Section IV and Section V, we propose the T-ELLM scheme and analyze its convergence properties. Afterward, we present the simulation results in Section VI. Finally, conclusions are stated in Section VII. In Table I, we summarize the main notations used in this paper.
@@ II System Model and Problem Formulation
TABLE I: List of key notations used in the paper
@@
$N$ Total number of devices
@@
$I$ The number of iterations for local training
@@ II-A1 Wireless FL Model:
As illustrated in Fig. 1, FL algorithms typically solve (2) through an iterative three-step process, which encompasses local training, aggregation, and synchronization, repeated over multiple rounds. Taking the example of FedAvg, in the $t$-th communication round, the central server broadcasts the $(t-1)$-th global model parameters $\bm{\omega}_{G}^{t-1}$ to each device, where $\bm{\omega}_{G}^{0}$ indicates an initialized model. After getting the global model, each device $n$ uses its local dataset to update the parameters of the local model $I$ times, i.e.,
@@ II-A2 Computation and Communication Model for FL:
  • Local Computation[31]: Let $f_{t,n}$ denote the CPU frequency (in CPU cycles per second) of device $n$ in the $t$-th round. The time required for $I$ times of local model training on device $n$ in the $t$-th round can be expressed as:
    $T^{\text{cmp}}_{t,n}=\frac{C|D_{n}|I}{f_{t,n}},$ (5)
    where $C$ (cycles per sample) represents the number of CPU cycles needed to process one sample by the backpropagation algorithm. Additionally, the energy consumption of device $n$ for local model training in the $t$-th round is given by
    $E^{\text{cmp}}_{t,n}=\zeta C|D_{n}|If^{2}_{t,n},$ (6)
    where $\zeta$ is the effective switched capacitance that depends on the chip architecture.
  • Model Transmission: After completing local model training, the selected devices upload their local model parameters to the BS via Frequency Division Multiple Access (FDMA). The achievable transmission rate of local device $n$ in the $t$-th round is given by
    $v_{t,n}=B_{t,n}\log_{2}\left(1+\frac{p_{t,n}G_{t,n}}{N_{0}B_{t,n}}\right),$ (7)
    where $B_{t,n}$ and $p_{t,n}$ denotes the bandwidth and transmission power allocated to device $n$ in the $t$-th communication round respectively, $G_{t,n}$ is the channel coefficient of device $n$ in the $t$-th round under the current environmental conditions, and $N_{0}$ is the noise power spectral density. Let $s_{n}$ denote the size (in bits) of the model parameters transmitted by device $n$. Therefore, the time required for device $n$ to upload its model parameters in the $t$-th communication round is given by
    $T^{\text{com}}_{t,n}=\frac{s_{n}}{v_{t,n}}.$ (8)
    Subsequently, the energy consumption of device $n$ in the $t$-th round for uploading model parameters can be expressed as:
    $E^{\text{com}}_{t,n}=p_{t,n}T^{\text{com}}_{t,n}=\frac{p_{t,k}s_{n}}{v_{t,n}}.$ (9)
@@ II-B Problem Formulation
A widely applied solver to the joint optimization problem $\mathfrak{P}1$ is DRL[27, 28, 16]. Generally, the dynamic process of the wireless FL can be formally modeled as the Markov Decision Process (MDP), denoted by $\mathcal{M}=(\mathcal{S},\mathcal{A},P,\mathcal{R},\gamma)$, where $\mathcal{S}$ is the state space with $\mathbf{s}_{t}\in\mathcal{S}$, and $\mathbf{s}_{t}$ can be composed of two parts, i.e.,
@@
where $\mathbb{R}_{t}$ represents the allocated resources, such as $f_{t,n}$, $p_{t,n}$, and $B_{t,n}$, to enable the full participation of devices within the subset $\mathbb{S}_{t}$, while we call such a subset as a feasible subset [^1] Notably, if $\mathbb{R}_{t}$ can only support the participation of a subset $\hat{\mathbb{S}}_{t}\subset\mathbb{S}_{t}$ of devices, we will denote the action as $[\hat{\mathbb{S}}_{t},\mathbb{R}_{t}]$ to avoid the ambiguity while the feasible subset is denoted as $\hat{\mathbb{S}}_{t}$.. $P$ is the transition function $P(\mathbf{s}_{t+1}|\mathbf{s}_{t},\mathbf{a}_{t})$. $\mathcal{R}$ is the reward function calculated by (12) and $\gamma$ is the discount factor. Generally, the NNs trained by DRL are only suitable for the current interactive environment, and thus, they need to be re-trained in a new environment. In addition, for different scenarios with distinctive resource constraints and task requirements, the NNs might be redesigned from scratch. For example, for two scenarios with different managed resources, the action space must be redesigned. Unfortunately, for a complicated problem that involves multiple interdependent decision-making tasks, the design and training of NNs becomes increasingly complex. Besides, since the number of neurons grows exponentially with the number of decision variables, such as $f_{n}$, $p_{n}$, and $B_{n}$ per device, it commonly encounters scalability issues. Consequently, an excessively large decision-making action space adversely affects the efficacy of DRL. Therefore, we propose to introduce LLM to address these challenges.

### v2

@@ Abstract
Federated Learning (FL) enables distributed model training across edge devices in a privacy-friendly manner. However, its efficiency heavily depends on effective device selection and high-dimensional resource allocation in dynamic and heterogeneous wireless environments. Conventional methods demand a confluence of domain-specific expertise, extensive hyperparameter tuning, and/or heavy interaction cost. This paper proposes a Tool-aided Evolutionary Large Language Model (T-ELLM) framework to generate a qualified policy for device selection in a wireless FL environment. Unlike conventional optimization methods, T-ELLM leverages natural language-based scenario prompts to enhance generalization across varying network conditions. The framework decouples the joint optimization problem mathematically, enabling tractable learning of device selection policies while delegating resource allocation to convex optimization tools. To facilitate the evolutionary process, T-ELLM interacts with a sample-efficient, model-based virtual learning environment that captures the relationship between device selection and learning performance. This developed virtual environment reduces reliance on real-world interactions, thus minimizing communication overhead while refining the LLM-based decision-making policy through group relative policy optimization. Theoretical analysis proves that the discrepancy between virtual and real environments is bounded, ensuring the advantage function learned in the virtual environment maintains a provably small deviation from real-world conditions. Experimental results demonstrate that T-ELLM outperforms benchmark methods in energy efficiency and exhibits robust adaptability to environmental changes.
@@ I Introduction
Next-Generation (xG) wireless communication systems are envisioned to support various intelligent applications and services[1], empowered by the exponential growth of wireless edge devices, such as mobile phones and sensors. As a prominent paradigm[2, 3], Federated Learning (FL) emerges by facilitating collaborative model training across decentralized devices while maintaining data locality in a privacy-friendly manner. Nevertheless, the deployment of FL in wireless networks faces severe challenges, arising from the underlying heterogeneous computation and communication capabilities across devices [4, 5]. Correspondingly, a proper resource management policy for FL, which selects the participating devices and calibrates the utilized communications and computing resources, will need to be developed in dynamic wireless environments [6]. Typically, such a problem results in a high-dimensional optimization problem that can be partially solved by heuristic solutions [7, 8] or Reinforcement Learning (RL)-based approaches [9, 10, 11, 12, 13]. However, heuristic solutions often demand a confluence of domain-specific expertise and extensive tuning before adapting to unseen scenarios, while RL-based approaches require heavy interactions with the environment and become sluggish for large system dimensions and changing system dynamics [14, 15, 16]. Therefore, there is a strong incentive to find alternative solutions[17].
@@ I-A Related Works
To implement FL over wireless networks, in each training round, edge devices upload the locally trained updates to a centralized server via wireless links, in exchange for aggregated models. Through several training rounds, the performance of the eventually learned global model is primarily impacted by resource and data heterogeneity. For example, disparities in computational and communication capabilities across devices lead to imbalanced time and energy consumption, while varying dataset sizes and non-Independent and Identically Distributed (non-IID) data distributions can result in biased model updates. Both would degrade the learning effectiveness [4, 5]. Additionally, the deployment of FL encounters constraints due to the energy and time [8] budget [10, 18]. For instance, the authors in [19, 20] formulate the joint learning and communication problem as the minimization of total energy consumption under latency constraints and provide computationally efficient closed-form solutions to optimize critical resources, including bandwidth allocation, CPU frequency, and transmission power. But, full device participation is bluntly assumed in [19, 20]. However, mobilizing the participation of a subset of qualified edge devices in each training round can contribute to improving learning efficiency. For example, biasing client selection towards clients with higher local loss [21] proves to yield faster convergence than the random selection in Federated Averaging (FedAvg) [2]. Also, a timer can be heuristically set to avoid the participation of straggler devices [8]. In [7], a guided participant selection scheme improves FL performance by jointly optimizing system and data utility through an epsilon-greedy algorithm. Moreover, a joint device scheduling and bandwidth allocation policy, which maximizes model accuracy within a fixed training time budget by a hybrid greedy-optimization approach, is proposed in [22], while energy constraints are neglected therein. More importantly, adapting these heuristics to unexpected scenarios, such as tuning hyperparameters to approximate FL performance limits[8, 22] or setting appropriate timers[7] when system bandwidth or device resources (e.g., maximum transmission power, dataset size) change, typically demands iterative adjustments and domain-specific expertise.
@@
RL provides an alternative methodology to solve wireless FL problems. Through continuously interacting with the environment, RL agents evaluate candidate policies in the wireless FL environment and gradually learn appropriate real-time policies. For example, [12] develops a Deep RL (DRL) method to select devices to upload their local model for aggregation, to maximize validation accuracy. Nevertheless, it overlooks energy and time constraints and assumes all devices participate in local training. Similarly, [11] adopts a Q-learning-based method to identify near-optimal participating devices. Furthermore, to ameliorate the impact of high dimensionality on single-agent RL with large-scale devices, [9] introduces a Multi-Agent Reinforcement Learning (MARL) framework that jointly optimizes model accuracy, processing latency, and communication efficiency. On the other hand, [13] improves the energy efficiency of FL by managing the CPU-cycle frequency of mobile devices based on DRL. To further improve the efficiency of FL, the works in [23, 24, 25] leverage DRL-based approaches to jointly optimize the device selection and bandwidth allocation. However, the above DRL-based methods only optimize partial parameters and rely on extensive training in a given environment. Therefore, the changes in environment or data often imply cumbersome retraining or even architectural redesign. Unfortunately, a real-world wireless scenario constantly evolves, and it inevitably induces substantial engineering cost. Meanwhile, direct interaction with the environment incurs substantial communication overhead. Although imitation learning [15] can acquire a policy by learning from the optimal trajectories to reduce interaction overhead, there are no such expert systems strictly in this FL scenario. Besides, the training difficulty, which increases exponentially along with the potentially involved devices and managed resources, limits the practicability [14, 16].
@@
The remarkable capabilities of transformer-based Large Language Models (LLMs) [26, 27, 28, 29] demonstrate the feasibility of a single general-purpose model, trained on extensive text corpora, for diverse, generalizable task accomplishment [30]. Capitalizing on the success of LLMs, we resort to an LLM-driven textual description approach to leverage its inherent flexibility and generalization. Leveraging fine-tuned LLMs for FL optimization starkly contrasts with those efforts, which focus on enhancing the training capabilities of LLMs in wireless FL [31, 32]. However, directly applying standard LLM methodologies to this specialized domain presents several fundamental obstacles. Prominently, common approaches like training and fine-tuning require vast real-world interaction datasets, which are prohibitively time-consuming and costly to collect [33, 34]. The non-availability of expert decision-making data for FL efficiency can add to the difficulty of reliably teaching an LLM to interpret wireless FL states and generate effective actions. Furthermore, general-purpose LLMs are inherently weak at precise mathematical reasoning and can be prone to hallucination[35], making them ill-suited for optimization-centric tasks, such as resource management in wireless FL. Meanwhile, the high latency associated with processing overlong token sequences for complex reasoning makes the sole adoption of LLM infeasible in time-sensitive wireless FL systems [36]. Therefore, due to the domain-specific expertise and computationally intensive optimization operations, extra effort is still needed to make LLMs qualified for decision-making in wireless FL.
TABLE I: The summary of differences with related literature on efficient FL decision-making.
Performance Guarantee Adaptability to Environments Artificial Design Complexity Interaction Cost Brief Description
[…]
@@ I-B Contributions
In this paper, we propose a Tool-aided Evolutionary LLM (T-ELLM) approach to realize efficient device selection and resource allocation in wireless FL. Specifically, T-ELLM adopts a natural language-based scenario prompt, where linguistic flexibility helps to improve the generalization capability[37]. Afterward, on top of a mathematics-established problem decoupling, T-ELLM capably yields device selection and resource management results by answering the scenario query and invoking convex optimization tools, respectively. Furthermore, to better tackle possible environmental changes, alongside available convex optimization tools, T-ELLM builds a sample-efficient model-based virtual learning environment, which characterizes the relationship between learning effectiveness and device selection. Once this high-fidelity model is established, the T-ELLM agent can be trained for a vast number of episodes entirely offline, at virtually zero communication cost. In other words, since the interaction cost is entirely induced by collecting samples for the simpler virtual environment modeling, this model-based approach yields high sample efficiency by drastically reducing the communication and time overhead for practically interacting LLMs with the real world. The main contributions of this paper are summarized as follows:
  • We consider a high-dimensional decision-making problem in wireless FL scenarios, and adopt a tool-based evolutionary LLM approach, which is referred to as T-ELLM. In particular, we fine-tune an LLM, which can respond to the linguistic description of wireless FL scenarios with an appropriate output of device selection. The astonishing contextual understanding capability in LLM empowers T-ELLM to cope with environmental changes flexibly, significantly reducing manual design effort and improving adaptability.
@@ I-C Paper Structure
The remainder of this paper is organized as follows. We introduce the system model and problem formulation in Section II. Afterward, Section III presents how to decouple the optimization problem and explains the rationale behind it. In Section IV and Section V, we propose the T-ELLM scheme and analyze its convergence properties. Afterward, we present the simulation results in Section VI. Finally, conclusions and future works are stated in Section VII.
@@ II System Model and Problem Formulation
In this section, we first present the FL model, followed by the definition of a wireless communication system model. Finally, we formulate the problem of efficient wireless FL resource management.
TABLE II: blueList of key notations used in the paper.
@@
NN Total number of devices
@@
II The number of iterations for local training
@@ II-A1 Wireless FL Model
As illustrated in Fig. 1, FL algorithms typically solve (2) through an iterative three-step process, which encompasses local training, aggregation, and synchronization, repeated over multiple rounds. Taking the example of FedAvg [2], in the $t$-th communication round, the central server broadcasts the (t1)(t-1)-th global model parameters $\bm{\omega}_{G}^{t-1}$ to each device, where $\bm{\omega}_{G}^{0}$ indicates an initialized model. After getting the global model, each device $n$ uses its local dataset to update the parameters of the local model $I$ times, i.e.,
@@ II-A2 Computation and Communication Model for FL:
  • Local Computation[23]: Let $f_{t,n}$ denote the CPU frequency (in CPU cycles per second) of device $n$ in the $t$-th round. The time required for $I$ times of local model training on device $n$ in the $t$-th round can be expressed as $T^{\text{cmp}}_{t,n}=\frac{C|D_{n}|I}{f_{t,n}}$, where $C$ (cycles per sample) represents the number of CPU cycles needed to process one sample by the backpropagation algorithm. Additionally, the energy consumption of device $n$ for local model training in the $t$-th round is given by $E^{\text{cmp}}_{t,n}=\zeta C|D_{n}|If^{2}_{t,n}$, where $\zeta$ is the effective switched capacitance that depends on the chip architecture.
  • Model Transmission: After completing local model training, the selected devices upload their local model parameters to the BS via Frequency Division Multiple Access (FDMA). The achievable transmission rate of local device $n$ in the $t$-th round is given by $v_{t,n}=B_{t,n}\log_{2}\left(1+\frac{p_{t,n}G_{t,n}}{N_{0}B_{t,n}}\right)$, where $B_{t,n}$ and $p_{t,n}$ denotes the bandwidth and transmission power allocated to device $n$ in the $t$-th communication round respectively, $G_{t,n}$ is the channel coefficient of device $n$ in the $t$-th round under the current environmental conditions, and $N_{0}$ is the noise power spectral density. Let $s_{n}$ denote the size (in bits) of the model parameters transmitted by device $n$. Therefore, the time required for device $n$ to upload its model parameters in the $t$-th communication round is given by $T^{\text{com}}_{t,n}=\frac{s_{n}}{v_{t,n}}$. Subsequently, the energy consumption of device $n$ in the $t$-th round for uploading model parameters can be expressed as $E^{\text{com}}_{t,n}=p_{t,n}T^{\text{com}}_{t,n}=\frac{p_{t,k}s_{n}}{v_{t,n}}$.
@@ II-B Problem Formulation
A widely applied solver to the joint optimization problem $\mathfrak{P}1$ is DRL[19, 20, 10]. Generally, the dynamic process of the wireless FL can be formally modeled as the Markov Decision Process (MDP), denoted by $\mathcal{M}=(\mathcal{S},\mathcal{A},P,R,\gamma)$, where $\mathcal{S}$ is the state space with $\mathbf{s}_{t}\in\mathcal{S}$, and $\mathbf{s}_{t}$ can be composed of two parts, i.e.,
@@
where $\mathbb{R}_{t}$ represents the allocated resources, such as $f_{t,n}$, $p_{t,n}$, and $B_{t,n}$, to enable the full participation of devices within the subset $\mathbb{S}_{t}$, while we call such a subset as a feasible subset [^1] Notably, if $\mathbb{R}_{t}$ can only support the participation of a subset $\hat{\mathbb{S}}_{t}\subset\mathbb{S}_{t}$ of devices, we will denote the action as $\hat{\mathbb{S}}_{t},\mathbb{R}_{t}]$ to avoid the ambiguity while the feasible subset is denoted as $\hat{\mathbb{S}}_{t}$.. $P$ is the transition function $P(\mathbf{s}_{t+1}|\mathbf{s}_{t},\mathbf{a}_{t})$. $R$ is the reward function calculated by (7) and $\gamma$ is the discount factor. Generally, the NNs trained by DRL are only suitable for the current interactive environment, and thus, they need to be re-trained in a new environment. In addition, for different scenarios with distinctive resource constraints and task requirements, the NNs might be redesigned from scratch. For example, for two scenarios with different managed resources, the action space must be redesigned. Unfortunately, for a complicated problem that involves multiple interdependent decision-making tasks, the design and training of NNs becomes increasingly complex. Besides, since the number of neurons grows exponentially with the number of decision variables, such as $f_{n}$, $p_{n}$, and $B_{n}$ per device, it commonly encounters scalability issues. Consequently, an excessively large decision-making action space adversely affects the efficacy of DRL. Therefore, we propose to introduce LLM to address these challenges.